# Online Graph Learning via Time-Vertex Adaptive Filters: From Theory to Cardiac Fibrillation

### v1

@@ Abstract
Graph Signal Processing (GSP) provides a powerful framework for analysing complex, interconnected systems by modelling data as signals on graphs. Recent advances in GSP have enabled the learning of graph structures from observed signals, but these methods often struggle with time-varying systems and real-time applications. Adaptive filtering techniques, while effective for online learning, have seen limited application in graph topology estimation from a GSP perspective. To this end, we introduce AdaCGP, an online algorithm for adaptive estimation of the Graph Shift Operator (GSO) from multivariate time series. The GSO is estimated from an adaptive time-vertex autoregressive model through recursive update formulae designed to address sparsity, shift-invariance and bias. Through simulations, we show that AdaCGP performs consistently well across various graph topologies, and achieves improvements in excess of 82% for GSO estimation compared to baseline adaptive vector autoregressive models. In addition, our online variable splitting approach for enforcing sparsity enables near-perfect precision in identifying causal connections while maintaining low false positive rates upon optimisation of the forecast error. Finally, AdaCGP’s ability to track changes in graph structure is demonstrated on recordings of ventricular fibrillation dynamics in response to an anti-arrhythmic drug. AdaCGP is shown to be able to identify the stability of critical conduction patterns that may be maintaining the arrhythmia in an intuitive way, together with its potential to support diagnosis and treatment strategies.
@@
Graph topology learning, adaptive graph signal processing, graph shift operator, time-vertex stochastic process, multivariate statistical models, cardiac fibrillation.
@@ I Introduction
In an era of unprecedented data generation, the challenge of understanding complex, time-varying systems has become increasingly critical. Much of this data arrives as simultaneous, long-running time series from diverse sources including financial markets, environmental monitoring stations [2], energy grid networks [3], and biomedical sensors [4]. While often appearing unstructured, these datasets frequently encode underlying spatiotemporal relationships that can reveal the underlying system dynamics.
@@
This is particularly evident in complex cardiac arrhythmias, such as Ventricular Fibrillation (VF) and Atrial Fibrillation (AF), where chaotic electrical propagation manifests as patient-specific ‘electrophenotypes’ [5, 6]. The ability to capture the electrophenotype and assess its stability holds crucial diagnostic value, yet current treatments remain xr due to our inability to characterise these patterns, leading to poor outcomes and high recurrence rates [5, 7]. The development of adaptive, data-driven approaches to address this challenge is therefore essential, both for advancing cardiac fibrillation treatment and for analysing other complex systems.
Networks/graphs provide a powerful low-dimensional framework for representing relationships between data sources [8], with nodes representing sensors and edges capturing their dependencies. When the graph structure is known, various tools from GSP theory and machine learning can analyse signals residing on the graph [9, 10]. However, in many real-world scenarios, the underlying graph structure is unknown and must be inferred from observed data. The GSO, which can take forms such as the adjacency matrix or Laplacian matrix, represents this structure by encoding the relationships between nodes. While existing approaches based on sparse inverse covariance estimation [11], signal smoothness [12, 13], and signal diffusion [14, 15, 16] have shown promise in learning GSOs, they typically focus on learning variants of the Laplacian or symmetric matrices, and neglect temporal dependencies. Also, many applications require learning more general forms of the GSO, such as the directed weighted adjacency matrix, that can capture asymmetric relationships and potential causal connections between nodes.
@@
In the context of time series data, vector autoregressive (VAR) models are commonly used to uncover a causal graph representation between the variables, as popularised by Granger [17]. More recently, GSP methods have been proposed to learn graph structures from multivariate time series. For instance, Segarra et al. [18] investigated the joint identification of graph filters and input signals, and Mei and Moura [19] proposed estimating GSOs through an autoregressive time-vertex system identification problem. However, these methods typically operate in a batch processing mode, which becomes inefficient and impractical when dealing with streaming data from complex systems where the graph structure is often time-varying [20]. In such cases, batch processing would require rerunning the model each time new data arrives, leading to high computational costs.
@@
Adaptive signal processing techniques are well-suited for such online learning scenarios. Algorithms like the Least Mean Square (LMS) [21], Recursive Least Squares (RLS) [22], and RLS-like LMS variants [23] provide powerful tools for learning and tracking time-varying parameters. Recent works have introduced online schemes for estimating causal graphs from sparse VAR processes, such as TISO (Topology Identification via Sparse Online learning) and its RLS counterpart TIRSO [24]. While adaptive filtering techniques have been extended to GSP [25, 26, 27], current research has focused on signal estimation and sampling on graphs. The problem of adaptive graph learning from streaming data remains largely unexplored.
To address this gap, we leverage time-vertex GSP theory to learn and track changes in weighted graph adjacency matrices. Our main contributions are:
@@
    An online time-vertex adaptive filtering algorithm, termed AdaCGP, that extends Mei and Moura’s [19] offline approach for graph structure learning;
@@
    A rigorous empirical analysis on synthetic data demonstrating the convergence and superiority of AdaCGP over adaptive VAR models, with prediction (forecast) error optimisation enabling reliable hyperparameter selection and GSO sparsity estimation;
@@ II Background
II-A Graph Structure and Signals
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a mathematical structure that captures the relationships between a set of entities. It consists of a set of $N$ nodes, $\mathcal{V}=\{v_{1},\dots,v_{N}\}$, and a set of edges, $\mathcal{E}$, representing their connections. A weight matrix, $\mathbf{W}\in\mathbb{R}^{N\times N}$, encodes the strength and pattern of the edge connections. This matrix, commonly referred to as the GSO in GSP, has entries $w_{ij}$ that are nonzero only if there is an edge connecting nodes $i$ and $j$, i.e., $(i,j)\in\mathcal{E}$. On this structure, a graph signal is defined as a function $f:\mathcal{V}\rightarrow\mathbb{R}$ that assigns a real value to each node, represented as a vector $\mathbf{x}=[x_{1},\dots,x_{N}]^{T}\in\mathbb{R}^{N\times 1}$, where $x_{n}$ is the signal value at node $v_{n}$.
@@ II-B Graph Filtering
$[H_{L}(\mathbf{W},\mathbf{h}),\mathbf{W}]=\mathbf{0},$ (2)
@@
$[H_{L}(\mathbf{W},\mathbf{h}_{1}),H_{K}(\mathbf{W},\mathbf{h}_{2})]=\mathbf{0}.$ (3)
Here, the commutator notation [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] for two operators $\mathcal{A}$ and $\mathcal{B}$ is defined as
$[\mathcal{A},\mathcal{B}]\triangleq\mathcal{A}\mathcal{B}-\mathcal{B}\mathcal{A},$ (4)
where $[\mathcal{A},\mathcal{B}]=0$ indicates that the operators commute [9].
@@ II-C Causal Graph Process
where $\mathbf{x}_{t}$ is the graph signal at time step $t$, $H_{p}(\mathbf{W},\mathbf{h}_{p})$ is an order $p$ graph filter as defined in (1), and $\mathbf{w}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N})$.
The CGP model is causal in the sense that the signal at a node at time $t$ can only be influenced by its $P$-hop neighbourhood from the previous $P$ time steps. Mathematically, this is reflected in the order of the polynomial coefficients being bounded by the lag $p$, $p=1,\ldots,P$.
This structure models information propagation through the graph at a fixed speed (one graph shift per sampling period). While more general formulations that decouple the polynomial order from the time lag are more expressive [28], the CGP model’s coupled temporal and spatial dependencies provide a tractable framework for graph learning [19]. The sampling frequency of discrete-time models can also be chosen to align with the propagation speed, making this assumption rather reasonable in practice.
@@
The CGP model is inherently weakly stationary due to its formulation as a sum of shift-invariant graph filters (as defined in [29]). While stationarity assumptions enable the definition of meaningful optimality criteria for filter design, real-world scenarios often involve non-stationary signals and time-varying graph structures. This motivates adaptive approaches that can track gradual changes through recursive optimisation of these criteria.
@@ III Objectives for Online Optimisation
where $\mathbf{h}=[\mathbf{h}_{1}^{T},...,\mathbf{h}_{P}^{T}]^{T}\in\mathbb{R}^{M\times 1}$ with $M={P(P+3)/2}$, $\mathbf{x}_{\tau}=\mathbf{0}$ for $\tau\leq 0$, and $\lambda\in(0,1]$. It is important to note that (6) represents a non-convex optimisation problem due to the polynomial in $\mathbf{W}$. To address this, we follow the approach proposed by Mei and Moura [19] and cast the problem into alternating steps of convex regularised least squares sub-problems.
@@ III-A Solving for $\boldsymbol{\Psi}_{p}=H_{p}(\mathbf{W},\mathbf{h}_{p})$
where $\boldsymbol{\Psi}=[\boldsymbol{\Psi}_{1},...,\boldsymbol{\Psi}_{P}]\in\mathbb{R}^{N\times NP}$ is a concatenation of the $P$ graph filters, vec()($\cdot$)( ⋅ ) is a vectorisation operator that stacks the columns of the matrix it acts on, and ${\|\cdot\|}_{1}$ is an $\ell_{1}$ norm, while $\mu_{p}$ is a constant which adjusts the degree of sparsity of the corresponding $\hat{\boldsymbol{\Psi}}_{p}$.
@@ III-C Estimating $\mathbf{h}$
Upon estimating $\hat{\mathbf{W}}$, we can re-frame the initial objective in (6) as a quadratic optimisation with respect to $\mathbf{h}$. Assuming sparsity in $\mathbf{h}$, we can substitute (1) into (6) and rearrange to give
@@ IV Online Estimation Algorithm
We introduce AdaCGP (Adaptive identification of Causal Graph Processes), an algorithm designed to achieve sparse online solutions to the objective in (6). While existing online methods such as $\ell_{1}$-regularised LMS [31] and oracle algorithms [32] achieve low MSE, they rarely produce truly sparse solutions compared to offline methods like basis pursuit [33]. This is an important limitation in our context for estimating $\hat{\mathbf{W}}$, since sparsity directly determines the presence or absence of edges in the graph structure, driving a causal interpretation.
To overcome this issue, we adapt the offline variable splitting approach of Schmidt et al. [34] to the online objective in (6), reformulating the alternating $\ell_{1}$-regularised sub-problems in (8) and (9) by splitting the target variables ($\hat{\boldsymbol{\Psi}}$ and $\hat{\mathbf{W}}$) into their positive and negative components as
@@
where $(\cdot)_{+}\geq\mathbf{0}$ and $(\cdot)_{-}\geq\mathbf{0}$ contain only the positive and negative parts of ()($\cdot$)( ⋅ ), respectively. The $\ell_{1}$-norm can then be expressed as a product-weighted sum. For the matrices $\hat{\boldsymbol{\Psi}}$ and $\hat{\mathbf{W}}$, this is given by
$\begin{split}&\|\hat{\boldsymbol{\Psi}}\|_{1}=\text{Tr}\left(\mathbf{1}_{N\times N}\hat{\boldsymbol{\Psi}}_{+}\right)+\text{Tr}\left(\mathbf{1}_{N\times N}\hat{\boldsymbol{\Psi}}_{-}\right),\\ &\|\hat{\mathbf{W}}|_{1}=\text{Tr}\left(\mathbf{1}_{N\times N}\hat{\mathbf{W}}_{+}\right)+\text{Tr}\left(\mathbf{1}_{N\times N}\hat{\mathbf{W}}_{-}\right)\\ \end{split}$
@@ IV-A Updating $\boldsymbol{\Psi}_{t}$
where we use $(\cdot)_{t}$ to denote variables ()($\cdot$)( ⋅ ) at the time instant $t$. We define $\hat{\boldsymbol{\Psi}}_{t}$, $\mathbf{M}_{t}$, and $\mathbf{Q}_{t}$ as the concatenation of graph filters, sparsity and commutative terms over all $P$, given by
@@
where it can be shown straightforwardly that $\nabla^{(t)}_{{(\hat{\boldsymbol{\Psi}}_{p})}}=\nabla^{(t)}_{{(\hat{\boldsymbol{\Psi}}_{p})_{+}}}$ and $\nabla^{(t)}_{{(\hat{\boldsymbol{\Psi}}_{p})_{-}}}=-\nabla^{(t)}_{{(\hat{\boldsymbol{\Psi}}_{p})}}$ for all $p$.
@@
where $\mathbf{A}_{t}=\text{diag}(\alpha_{1},\alpha_{2},...,\alpha_{P})\in\mathbb{R}^{P\times P}$ is a diagonal matrix of stepsizes for each filter block. Similarly, the update equation for ${\hat{\boldsymbol{\Psi}}}_{-_{t}}$ can be obtained as

### v2

# Online Graph Topology Learning via Time-Vertex Adaptive Filters: From Theory to Cardiac Fibrillation

@@ Abstract
Graph Signal Processing (GSP) provides a powerful framework for analysing complex, interconnected systems by modelling data as signals on graphs. While recent advances have enabled graph topology learning from observed signals, existing methods often struggle with time-varying systems and real-time applications. To address this gap, we introduce AdaCGP, a sparsity-aware adaptive algorithm for dynamic graph topology estimation from multivariate time series. AdaCGP estimates the Graph Shift Operator (GSO) through recursive update formulae designed to address sparsity, shift-invariance, and bias. Through comprehensive simulations, we demonstrate that AdaCGP consistently outperforms multiple baselines across diverse graph topologies, achieving improvements exceeding 83% in GSO estimation compared to state-of-the-art methods while maintaining favourable computational scaling properties. Our variable splitting approach enables reliable identification of causal connections with near-zero false alarm rates and minimal missed edges. Applied to cardiac fibrillation recordings, AdaCGP tracks dynamic changes in propagation patterns more effectively than established methods like Granger causality, capturing temporal variations in graph topology that static approaches miss. The algorithm successfully identifies stability characteristics in conduction patterns that may maintain arrhythmias, demonstrating potential for clinical applications in diagnosis and treatment of complex biomedical systems.
@@
Graph topology estimation, functional connectivity, adaptive graph signal processing, graph shift operator, time-vertex stochastic process, multivariate statistical models, cardiac fibrillation.
@@ I Introduction
In an era of unprecedented data generation, the challenge of understanding complex, time-varying systems has become increasingly critical. Much of this data arrives as simultaneous, long-running time series from diverse sources including financial markets, environmental monitoring stations [2], energy grid networks [3], and biomedical sensors [4]. While often appearing unstructured, these datasets frequently encode underlying spatio-temporal relationships that can reveal system dynamics and topology.
Networks/graphs provide a powerful low-dimensional framework for representing relationships between data sources [5], with nodes representing sensors and edges capturing their dependencies. When the graph topology (the GSO) is known, various tools from graph signal processing (GSP) theory can analyse signals residing on the graph [6, 7]. However, in many real-world scenarios, the underlying graph topology is unknown and must be estimated from observed data, often in an online/adaptive manner as new measurements become available.
@@
This fundamental challenge, known as graph topology estimation, has attracted significant attention in recent years, particularly in clinical medicine, where graph topology estimates may reflect functional connectivity in biomedical systems. For instance, in complex cardiac arrhythmias such as Ventricular Fibrillation (VF) and Atrial Fibrillation (AF), chaotic electrical propagation manifests as patient-specific ‘electrophenotypes’ [8, 9]. The ability to capture these electrophenotypes and assess their stability holds crucial diagnostic value, yet current treatments remain ineffective due to our inability to characterise these patterns, leading to poor outcomes and high recurrence rates [8, 10]. Recent work has applied graph topology estimation to identify fibrillation electrophenotypes [11], but these offline/batch approaches lack the ability to track evolving network dynamics and measure the electrophenotype’s stability.
Beyond cardiac applications, many neurological and psychiatric conditions, including epilepsy, Alzheimer’s disease, and schizophrenia, have been increasingly recognised as disorders of dysregulated network connectivity [12]. Graph-based functional connectivity characterises the interactions between brain regions in both health and disease [12]. Conventional approaches such as Pearson’s correlation provide a simplistic representation of functional interactions [13], while precision matrix methods offer a more refined modelling of indirect influences [14, 15]. While dynamic functional connectivity techniques, such as employing batch methods in sliding windows, have improved the capture of temporal variations [16, 17], truly adaptive approaches for the task remain in their infancy.
Alongside neurological and cardiac applications, modelling complex drug-disease interactions represents another critical domain where graph topology estimation proves invaluable. Traditional methods for analysing these interactions often rely on pairwise statistical associations, which fail to capture the underlying causal mechanisms and network-level dependencies that drive disease progression and treatment response [18]. Graph-based models provide an effective alternative by representing biomedical systems as structured networks, where nodes correspond to biological entities (genes, proteins, drugs, physiological signals) and edges encode their functional relationships [19], while causal inference techniques, such as those popularised by Granger [20], allow the identification of directed interactions by testing whether past observations of one signal improve the prediction of another [21].
@@
Real-world applications have driven the development of diverse methodological approaches to graph topology estimation, spanning multiple analytical frameworks. Traditional approaches to graph topology estimation include methods based on sparse inverse covariance estimation [15], and GSP-based signal smoothness [22, 23] and diffusion [24, 25, 26]. While promising, these methods typically focus on learning variants of the Laplacian or symmetric matrices, neglecting temporal dependencies and directed relationships. For time series data, vector autoregressive (VAR) models have been widely used to uncover causal graph topologies [20]. More recently, GSP methods have extended these approaches, with [27] investigating joint identification of graph filters and input signals, and [28] proposing GSO estimation through autoregressive time-vertex system identification. Graph-based deep learning approaches have further advanced the field [29, 30]. However, a critical limitation of these methods is that they typically operate in batch processing mode, which becomes inefficient and impractical when dealing with streaming data from complex systems where the graph topology may be time-varying [31]. In such cases, batch processing would require rerunning the model each time new data arrives, leading to high computational costs.
@@
To overcome these fundamental limitations of batch processing approaches, adaptive signal processing techniques offer a promising alternative for online learning scenarios. Algorithms like the Least Mean Square (LMS) [32], Recursive Least Squares (RLS) [33], and RLS-like LMS variants [34] provide powerful tools for learning and tracking time-varying parameters. Recent works have introduced online schemes for estimating causal graphs from sparse VAR processes, such as TISO (Topology Identification via Sparse Online learning) and its RLS counterpart TIRSO [35], and sparse Structural Equation Model (SEM) processes [36]. However, achieving true sparsity in online settings remains challenging, as existing methods often struggle to explicitly zero out non-causal elements despite employing sparsity-promoting regularisation [37]. While adaptive filtering techniques have been extended to GSP [38, 39, 40], current research has focused primarily on signal estimation and sampling on graphs. The problem of adaptive graph topology estimation from streaming data remains largely unexplored, particularly sparsity-aware approaches that can reliably identify the presence or absence of causal connections.
To address this gap, we propose AdaCGP (Adaptive identification of Causal Graph Processes), which leverages time-vertex GSP theory to learn and track changes in weighted graph adjacency matrices through a novel online sparsity-aware topology estimation method. Our approach extends the offline approach for graph topology estimation in [28] to the online setting, enabling efficient processing of streaming data and tracking of time-varying graph topologies. Our main contributions are:
@@
    An online time-vertex adaptive filtering algorithm that extends offline approaches for graph topology estimation to streaming data scenarios, with a focus on accurate topology estimation;
@@
    A rigorous empirical analysis on synthetic data demonstrating the convergence and superiority of AdaCGP over adaptive VAR models, with prediction (forecast) error optimisation enabling reliable hyper-parameter selection and GSO sparsity estimation;
@@ II Background
II-A Graph Topology and Signals
A graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ is a mathematical structure that captures the relationships between a set of entities. It consists of a set of $N$ nodes, $\mathcal{V}=\{v_{1},\dots,v_{N}\}$, and a set of edges, $\mathcal{E}$, representing their connections. A weight matrix, $\mathbf{W}\in\mathbb{R}^{N\times N}$, encodes the strength and pattern of the edge connections. This matrix, commonly referred to as the GSO in GSP, has entries $w_{ij}$ that are non-zero only if there is an edge connecting nodes $i$ and $j$, i.e., $(i,j)\in\mathcal{E}$. On this structure, a graph signal is defined as a function $f:\mathcal{V}\rightarrow\mathbb{R}$ that assigns a real value to each node, represented as a vector $\mathbf{x}=[x_{1},\dots,x_{N}]^{T}\in\mathbb{R}^{N\times 1}$, where $x_{n}$ is the signal value at node $v_{n}$.
@@ II-B Graph Filtering
$H_{L}(\mathbf{W},\mathbf{h}),\mathbf{W}]=\mathbf{0},$ (2)
@@
$H_{L}(\mathbf{W},\mathbf{h}_{1}),H_{K}(\mathbf{W},\mathbf{h}_{2})]=\mathbf{0}.$ (3)
Here, the commutator notation $\cdot$$\cdot$] for two operators $\mathcal{A}$ and $\mathcal{B}$ is defined as
$\mathcal{A},\mathcal{B}]\triangleq\mathcal{A}\mathcal{B}-\mathcal{B}\mathcal{A},$ (4)
where $\mathcal{A},\mathcal{B}]=0$ indicates that the operators commute [6].
@@ II-C Causal Graph Process
where $\mathbf{x}_{t}$ is the graph signal at time step $t$, $H_{p}(\mathbf{W},\mathbf{h}_{p})$ is an order $p$ graph filter as defined in (1), and $\mathbf{w}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{N\times N})$.
The CGP model is causal in the sense that the signal at a node at time $t$ can only be influenced by its $P$-hop neighbourhood from the previous $P$ time steps. Mathematically, this is reflected in the order of the polynomial coefficients being bounded by the lag $p$, where $p=1,\ldots,P$.
This structure models information propagation through the graph at a fixed speed (one graph shift per sampling period). While more general formulations that decouple the polynomial order from the time lag are more expressive [41], the CGP model’s coupled temporal and spatial dependencies provide a tractable framework for graph topology estimation [28]. The sampling frequency of discrete-time models can also be chosen to align with the propagation speed, making this assumption rather reasonable in practice.
@@
The CGP model is inherently weakly stationary due to its formulation as a sum of shift-invariant graph filters (as defined in [42]). While stationarity assumptions enable the definition of meaningful optimality criteria for filter design, real-world scenarios often involve non-stationary signals and time-varying graph topologies. This motivates adaptive approaches that can track gradual changes through recursive optimisation of these criteria.
@@ III Objectives for Online Optimisation
where $\mathbf{h}=[\mathbf{h}_{1}^{T},...,\mathbf{h}_{P}^{T}]^{T}\in\mathbb{R}^{M\times 1}$ with M=P(P+3)/2M={P(P+3)/2}, $\mathbf{x}_{\tau}=\mathbf{0}$ for $\tau\leq 0$, and $\lambda\in(0,1]$. It is important to note that (6) represents a non-convex optimisation problem due to the polynomial in $\mathbf{W}$. To address this, we follow the approach proposed in [28] and cast the problem into alternating steps of convex regularised least squares sub-problems.
@@ III-A Solving for $\boldsymbol{\Psi}_{p}=H_{p}(\mathbf{W},\mathbf{h}_{p})$
where $\boldsymbol{\Psi}=[\boldsymbol{\Psi}_{1},...,\boldsymbol{\Psi}_{P}]\in\mathbb{R}^{N\times NP}$ is a concatenation of the $P$ graph filters, vec()($\cdot$) is a vectorisation operator that stacks the columns of the matrix it acts on, and ${\|\cdot\|}_{1}$ is an $\ell_{1}$ norm, while $\mu_{p}$ is a constant which adjusts the degree of sparsity of the corresponding $\hat{\boldsymbol{\Psi}}_{p}$.
@@ III-C Estimating $\mathbf{h}$
Upon estimating $\mathbf{W}$, we can re-frame the initial objective in (6) as a quadratic optimisation with respect to $\mathbf{h}$. Assuming sparsity in $\mathbf{h}$, we can substitute (1) into (6) and rearrange to give
@@ IV Online Estimation Algorithm
We introduce AdaCGP, an algorithm designed to achieve sparse online solutions to the objective in (6). While existing online methods such as $\ell_{1}$-regularised LMS [44] and oracle algorithms [45] achieve low MSE, they rarely produce truly sparse solutions compared to offline methods like basis pursuit [37]. This is an important limitation in our context for estimating $\mathbf{W}$, since sparsity directly determines the presence or absence of edges in the graph topology, driving a causal interpretation.
To overcome this issue, we adapt the offline variable splitting approach of [46] to the online objective in (6), reformulating the alternating $\ell_{1}$-regularised sub-problems in (8) and (9) by splitting the target variables ($\hat{\boldsymbol{\Psi}}$ and $\hat{\mathbf{W}}$) into their positive and negative components as
@@
where $(\cdot)_{+}\geq\mathbf{0}$ and $(\cdot)_{-}\geq\mathbf{0}$ contain only the positive and negative parts of $(\cdot)$, respectively. The $\ell_{1}$-norm can then be expressed as a product-weighted sum. For the matrices $\hat{\boldsymbol{\Psi}}$ and $\hat{\mathbf{W}}$, this is given by
[…]
@@ IV-A Updating $\boldsymbol{\Psi}_{t}$
where we use $(\cdot)_{t}$ to denote variables $(\cdot)$ at the time instant $t$. We define $\hat{\boldsymbol{\Psi}}_{t}$, $\mathbf{M}_{t}$, and $\mathbf{Q}_{t}$ as the concatenation of graph filters, sparsity and commutative terms over all $P$, given by
@@
where it can be shown straightforwardly that $\nabla^{(t)}_{{(\hat{\boldsymbol{\Psi}}_{p})_{+}}}=\nabla^{(t)}_{{(\hat{\boldsymbol{\Psi}}_{p})}}$ and $\nabla^{(t)}_{{(\hat{\boldsymbol{\Psi}}_{p})_{-}}}=-\nabla^{(t)}_{{(\hat{\boldsymbol{\Psi}}_{p})}}$ for all $p$.
@@
where $\mathbf{A}_{t}=\text{diag}(\alpha_{1},\alpha_{2},...,\alpha_{P})\in\mathbb{R}^{P\times P}$ is a diagonal matrix of step-sizes for each filter block. Similarly, the update equation for ${\hat{\boldsymbol{\Psi}}}_{-_{t}}$ can be obtained as